OpenStudy (yas123):

write an equation in slope intercept form for the line containing the indicated points. (-1,-4) and (-2,5) i need help on that

OpenStudy (anonymous):

slope intercept form is (y-y1)=m(x-x1) where (x1,y1) is a point on the line (any point will work) and m is the slope. first we figure out the slope by doing y1-y2/x1-x2 using our two given points. so our slope is -4-5/-1--2 which is -9/1 =-9. we plug that into the equation along with the second point (either one works) y-5=-9(x--2) which is y-5=-9(x+2) which is y-5=-9x-18 and then solving for y gives us: y=-9x-13

OpenStudy (chaise):

m = y1-y2/x1-x2 m = -4-5/-1--2 m = 9/1 y=mx+b 5=9/1x-2 +b 5 = -18 +b 23 = b y = 9x +23 REMEMBER A NEGATIVE AND NEGATIVE MAKES POSITIVE.

OpenStudy (anonymous):

^ wrong chaise. -4-5 is definitely -9 there buddy. and yes -1--2 is a positive 1, so you have -9/1, which is what i have. thanks tho

OpenStudy (yas123):

wait so what is the correct answer b/s fishcx and chaise gave mi two diffrent answer

OpenStudy (anonymous):

my answer is correct. you can check it here: http://www.webmath.com/equline1.html